# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v1

@@ 6 Future Directions
LLMs can be conceptualized as parametric knowledge stores, where the parameters of the model—typically the weights of the neural network—encode a representation of the world’s knowledge. This parametric approach to knowledge storage means that the knowledge is implicitly embedded within the model’s architecture, and it can be retrieved and manipulated through the computational processes of the neural network Allen-Zhu and Li (2023b). In contrast, non-parametric knowledge storage involves methods where the knowledge is explicitly represented and can be directly accessed. Examples of non-parametric knowledge storage include knowledge graphs, databases, and symbolic reasoning systems, where knowledge is represented as discrete symbols or facts. Parametric knowledge enables LLMs to deeply compress and integrate information Huang et al. (2024c); Shwartz-Ziv and LeCun (2024), allowing them to generalize and apply this knowledge across various contexts. This is akin to LLMs mastering the mathematical operation rule of “mod” through parametric knowledge, enabling them to generalize and seamlessly solve all mod-related problems Pearce et al. (2023); Hu et al. (2024). Conversely, non-parametric knowledge requires extensive searches across the knowledge space for each user query. Subsequently, Wang et al. (2024a) also prove that non-parametric knowledge severely fails in complex reasoning tasks, with accuracy levels approaching random guessing. Unfortunately, parametric knowledge within LLMs is opaque, often encountering challenges such as interpretability issues, outdated information, hallucinations, and security concerns.
@@
Generally, inspired by the knowledge mechanisms analysis in LLMs, we have the potential to develop more architecture and learning strategy for organizing knowledge within LLMs. These efficient LLMs Sastry et al. (2024) are advancing toward lower GPU, computation, and storage resource requirements, as well as smaller model sizes by combining the strengths of parametric and non-parametric knowledge Yang et al. (2024b); Momeni et al. (2024); Chen (2024); Pan et al. (2024b, 2023b).
@@ 6.2 Embodied Intelligence
As discussed in the previous part, current language models are mainly based on tuning to gain knowledge, which requires a lot of training and high-quality data. These learnings are passive whereas, to be a human, evolution usually also undergoes communication and interaction. As an intelligent agent, the models should be able to learn through interactions and learn by themselves spontaneously. Recently, some work has attempted to enable the model to learn by themselves Zhang et al. (2024b) or learn by interaction with the environment Xu et al. (2024a); Xi et al. (2024). By integrating self-evolving mechanisms, models can continuously update their knowledge base and improve their understanding without relying solely on manually curated datasets. This not only reduces the dependency on large-scale labeled data but also allows the models to adapt to evolving linguistic norms and cultural contexts over time.
@@ 6.3 Domain LLMs
6.4 Trustworthy LLMs
Wei et al. (2023) delve into the safety of LLM and reveal that the success of jailbreak is mainly due to the distribution discrepancies between malicious attacks and training data. Geva et al. (2022b) and Wang et al. (2024b) further discover that some parameters within LLMs, called toxic regions, are intrinsically tied to the generation of toxic content. Ji et al. (2024a) even conjecture that LLMs resist alignment. Therefore, traditional aligned methods, DPO Rafailov et al. (2023) and SFT, seem to merely bypass toxic regions Lee et al. (2024); Wang et al. (2024b), making them susceptible to other jailbreak attacks Zhang et al. (2023d).
Inspired by the knowledge mechanism in LLMs, a promising trustworthy strategy may be designing architecture and training process during the pre-training phase to encourage monosemanticity and sparsity, which make the reverse engineering process more tractable Jermyn et al. (2022); Bricken et al. (2023); Liu et al. (2024b); Zhang et al. (2021); Tamkin et al. (2023). Yet, maintaining sparsity for a vast amount of world knowledge requires substantial resources, and whether monosemantic architecture can support advanced intelligence remains elusive. Besides, machine unlearning Nguyen et al. (2022); Tian et al. (2024) aims to forget privacy or toxic information learned by LLMs. However, these unlearning methods suffer over-fitting, forgetting something valuable. Another alternative technique is knowledge editing, precisely modifying LLMs using few instances during the post-training stage Mazzia et al. (2023); Yao et al. (2023); Wang et al. (2023d); Hase et al. (2024); Qian et al. (2024b). Extensive experiments demonstrate that knowledge editing has the potential to detoxify LLMs Yan et al. (2024). Specifically, Wu et al. (2023) and Geva et al. (2022b) deactivate the neurons related to privacy information and toxic tokens, respectively. However, knowledge editing also introduces side effects, such as the inability of the modified knowledge to generalize to multi-hop tasks Zhong et al. (2023); Li et al. (2023c); Cohen et al. (2023); Kong et al. (2024) and the potential to impair the model’s general capabilities Gu et al. (2024). Therefore, recent efforts aim to manipulate intermediate representation of LLMs during inference stage instead of editing parameters of LLMs Zou et al. (2023); Turner et al. (2023); Zhou et al. (2023b); Zhu et al. (2024). These representations can track and address a wide range of safety-relevant problems, including honesty, harmlessness, and power-seeking. Later, Wu et al. (2024) fine-tune representation for updating new knowledge in LLMs. However, fine-tuning representation requires meticulous hyperparameter adjustment for each task separately, necessitating efficient hyperparameter-tuning strategies that are compute- or time-matched.
@@ Appendix A Comparison of Methods for Knowledge Evolution
  • Machine Unlearning Nguyen et al. (2022); Tian et al. (2024); Liu et al. (2024c) focuses on discarding undesirable behaviors from LLMs.
@@ Appendix B Universality Intelligence
To validate the hypotheses in this paper across different architectures, we first introduce other popular model architectures in §B.1, and observe the generalizability of our hypotheses across other model architectures in §B.2. Besides, recent work further claims that models trained with different data, modalities, and objectives are converging to shared representation spaces Huh et al. (2024). Artificial and biological neural networks  [^8] Unless otherwise specified as biological networks, the terms models, neural networks, neural models, and machines refer to artificial neural networks. also share similar features and circuits, suggesting a universal underlying mechanism Sucholutsky et al. (2023); Chan et al. (2023); Kornblith et al. (2019). Therefore, analogous to biological taxonomy, we introduce artificial neural model family and and discuss the potential universality intelligence in the future in §B.3.
@@ B.1.3 Vision and Multi-modal Models
In the realm of vision and multi-modal models, various architectures have emerged, each with its unique approach to tackling complex visual tasks. For example, GANs (Generative Adversarial Nets) Goodfellow et al. (2014) consist of two neural networks: a generator and a discriminator. Through adversarial learning, the generator aims to produce realistic data samples (such as images), while the discriminator attempts to distinguish between real and generated data. Diffusion Model Li et al. (2023b); Sohl-Dickstein et al. (2015) is a powerful tool for generating high-quality images and data. It simulates a diffusion process by gradually adding and removing noise to achieve data generation. ResNet (Residual Network) He et al. (2016) introduced residual learning, revolutionizing deep network training by improving efficiency and performance through skip connections. ViT (Vision Transformer) Dosovitskiy et al. (2021) integrated the Transformer architecture into vision tasks, capturing long-range dependencies by processing image patches.
@@ B.3 Machine and Human
  • Species: BERT, GPT, Llama, Mistral, Mamba, etc., likened to “Sapiens”, “Pan troglodytes”, ‘Pan paniscus”.
Metaphorically, Llama-7B, Llama-13B, Llama-70B, etc., can be viewed as the infancy, childhood, adulthood of humans. Shah et al. (2024) further find that, regardless of model size, the developmental trajectories of PLMs consistently exhibit a window of maximal alignment with human cognitive development. Therefore, we hypothesize that artificial neural networks (machine) and biological neural networks (human) tend to converge to universality intelligence. In other words, huamn and machine share similar features and circuits.

### v2

@@ 7 Future Directions
LLMs can be conceptualized as parametric knowledge stores, where the parameters of the model—typically the weights of the neural network—encode a representation of the world’s knowledge. This parametric approach to knowledge storage means that the knowledge is implicitly embedded within the model’s architecture, and it can be retrieved and manipulated through the computational processes of the neural network Allen-Zhu and Li (2023b). In contrast, non-parametric knowledge storage involves methods where the knowledge is explicitly represented and can be directly accessed. Examples of non-parametric knowledge storage include knowledge graphs, databases, and symbolic reasoning systems, where knowledge is represented as discrete symbols or facts. Parametric knowledge enables LLMs to deeply compress and integrate information Huang et al. (2024d); Shwartz-Ziv and LeCun (2024), allowing them to generalize and apply this knowledge across various contexts. This is akin to LLMs mastering the mathematical operation rule of “mod” through parametric knowledge, enabling them to generalize and seamlessly solve all mod-related problems Pearce et al. (2023); Hu et al. (2024). Conversely, non-parametric knowledge requires extensive searches across the knowledge space for each user query. Subsequently, Wang et al. (2024a) also prove that non-parametric knowledge severely fails in complex reasoning tasks, with accuracy levels approaching random guessing. Unfortunately, parametric knowledge within LLMs is opaque, often encountering challenges such as interpretability issues, outdated information, hallucinations, and security concerns.
@@
Generally, inspired by the knowledge mechanisms analysis in LLMs, we have the potential to develop more architectural and learning strategies for organizing knowledge within LLMs. These efficient LLMs Sastry et al. (2024) are advancing toward lower GPU, computation, and storage resource requirements, as well as smaller model sizes by combining the strengths of parametric and non-parametric knowledge Yang et al. (2024b); Momeni et al. (2024); Chen (2024); Pan et al. (2024b, 2023b).
@@ 7.2 Embodied Intelligence
As discussed in the previous part, current language models are mainly based on tuning to gain knowledge, which requires a lot of training and high-quality data. These learnings are passive whereas, to be a human, evolution usually also undergoes communication and interaction. As an intelligent agent, the models should be able to learn through interactions and learn by themselves spontaneously. Recently, some work has attempted to enable the model to learn by themselves Zhang et al. (2024c) or learn by interaction with the environment Xu et al. (2024a); Xi et al. (2024). By integrating self-evolving mechanisms, models can continuously update their knowledge base and improve their understanding without relying solely on manually curated datasets. This not only reduces the dependency on large-scale labeled data but also allows the models to adapt to evolving linguistic norms and cultural contexts over time.
@@ Appendix A Comparison of Methods for Knowledge Evolution
  • Machine Unlearning Nguyen et al. (2022); Tian et al. (2024); Liu et al. (2024d) focuses on discarding undesirable behaviors from LLMs.
@@ Appendix B Universality Intelligence
To validate the hypotheses in this paper across different architectures, we first introduce other popular model architectures in §B.1, and observe the generalizability of our hypotheses across other model architectures in §B.2. Besides, recent work further claims that models trained with different data, modalities, and objectives are converging to shared representation spaces Huh et al. (2024). Artificial and biological neural networks  [^10] Unless otherwise specified as biological networks, the terms models, neural networks, neural models, and machines refer to artificial neural networks. also share similar features and circuits, suggesting a universal underlying mechanism Sucholutsky et al. (2023); Chan et al. (2023); Kornblith et al. (2019). Therefore, analogous to biological taxonomy, we introduce artificial neural model family and discuss the potential universality intelligence in the future in §B.3.
@@ B.1.3 Vision and Multi-modal Models
In the realm of vision and multi-modal models, various architectures have emerged, each with its unique approach to tackling complex visual tasks. For example, GANs (Generative Adversarial Nets) Goodfellow et al. (2014) consist of two neural networks: a generator and a discriminator. Through adversarial learning, the generator aims to produce realistic data samples (such as images), while the discriminator attempts to distinguish between real and generated data. Diffusion Model Li et al. (2023c); Sohl-Dickstein et al. (2015) is a powerful tool for generating high-quality images and data. It simulates a diffusion process by gradually adding and removing noise to achieve data generation. ResNet (Residual Network) He et al. (2016) introduced residual learning, revolutionizing deep network training by improving efficiency and performance through skip connections. ViT (Vision Transformer) Dosovitskiy et al. (2021) integrated the Transformer architecture into vision tasks, capturing long-range dependencies by processing image patches.
@@ B.3 Machine and Human
  • Species: BERT, GPT, Llama, Mistral, Mamba, etc., likened to “Sapiens”, “Pan troglodytes”, and “Pan paniscus”.
Metaphorically, Llama-7B, Llama-13B, Llama-70B, etc., can be viewed as the infancy, childhood, and adulthood of humans. Shah et al. (2024) further find that, regardless of model size, the developmental trajectories of PLMs consistently exhibit a window of maximal alignment with human cognitive development. Therefore, we hypothesize that artificial neural networks (machine) and biological neural networks (human) tend to converge to universality intelligence. In other words, huamn and machine share similar features and circuits.